### v1

@@ Abstract
[…]
@@ 1.1 Historical Background: The End of the Second Sino-Japanese War
We started narrating the historical background from 1943. On December 1st, 1943, the heads of state of the United States, China, and the United Kingdom jointly issued the Cairo Declaration, which stipulated the ownership of Taiwan after World War II: "It is their purpose that Japan shall be stripped of all the islands in the Pacific which she has seized or occupied since the beginning of the first World War in 1914, and that all the territories Japan has stolen from the Chinese, such as Manchuria, Formosa [^1] Formosa is another name for Taiwan., and The Pescadores, shall be restored to the Republic of China."[1]Similarly, the Potsdam Proclamation also stipulated that Taiwan belonged to China after the war.
On August 15th, 1945, Japan surrendered unconditionally, and the sovereignty of Taiwan was legally returned to China. On September 1st of the same year, the Chinese Nanjing National Government established the Taiwan Provincial Garrison Command in Chongqing. On September 3rd, Chiang Kai-shek issued an order: "Appoint General Chen Yi as the chief official for accepting the surrender of Taiwan and the Penghu Islands."[2]At 10 a.m. on October 10th, all officers and soldiers of the Forward Command Post, along with local gentry and civilians from Taiwan, gathered at the former Guild Hall [^2] It’s now known as Taipei Zhongshan Hall.in Taipeito hold the first National Day commemoration in Taiwan Province, declaring the return of Taiwan’s sovereignty.
@@
In August 1945, the price of rice once dropped below 2 Old Taiwan dollar(OTD [^3] Old Taiwan dollar (1946–1949), predecessor to today’s New Taiwan dollar, withdrawn at 40,000:1 owing to post-war hyperinflation. We abbreviate the Old Taiwan dollar as OTD.) per jin [^4] Jin is a traditional chinese unit of measurement. Commonly, one jin is equal to 500 grams.. From mid-August to mid-September 1945, the market in Taiwan Province once presented a scene of abundant goods and a steady decline in prices. However, From the second half of September in the 34th year of the Republic of China [^5] September 1945., the price of grain soared. In early 1947, the price of rice in Taipei rose to 14 OTD per jin. Later, under the influence of the gold rush in Shanghai, the price of rice suddenly jumped to over 30 OTD per jin, and even reached more than 40 OTD per jin on the black market.
@@ 2 Literature Review
Some scholars have conducted research on the food shortage during this period. Many scholars have studied the rice price data at some specific time points. Zeng Lei-lei pointed out that the price of rice in Taiwwan rose above 2 OTD per jin on October 10th and soared to 8 OTD/jin by the end of December[4]. Weng Jia-hsi: "In January 1946, one jin of rice cost only 8.84 OTD."[5]On page 265 of Weng’s paper, Weng presented a table of the prices of major daily necessities in Taipei City in the early post-war period. Among them, the price of rice was 8.84 OTD/jin in January 1946 and rose to 42.67 OTD/jin in February 1947, with an increase multiple of 4.83[5]. Professor Wang Xiaobo from Taiwan believes that according to the investigation data of some scholars, the price of rice in Taiwan was 6.3 OTD per jin in January 1946 and 32.33 OTD per jin in February 1947. However, we remain skeptical about the data provided by Wang Xiaobo because the gap between the data he provided and those provided by other scholars exceeds 23 percents, and it is even larger than the official data of the Taiwan government. We fully doubted that the data he provided are for other types of rice, not for top-quality rice.
Additionally, we have found many official documentary books that record the food shortage events that occurred in Taiwan during this period. The book "Taiwan’s Recovery and the Provincial Conditions of the Five Years After Recovery", edited by Chen Mingzhong and Chen Xingtang and published by Nanjing Publishing House, includes many reports from newspapers and periodicals during the grain shortage incident. We believe that the representative newspaper of these newspaper types is Wenhui Daily. These two individuals are researchers at the Second National Historical Archives of China, as we believe the content of this book is highly credible[6]. Given that this material has not been reprinted since 2008 and it cannot be found in the bookstore, to Science Data Bank[7] we uploaded the scanned copy whose filename is RefBook1.pdf [^6] DOI: 10.57760/sciencedb.32762. In March 1947, several reports in Wenhui Daily mentioned the price of rice in the Taiwanese market at that time. Among them, the reports of the two journalists caught our attention. Journalist Feng Yan’s report mentioned that the price of rice was 13.5 OTD per jin in January 1947, soon rose to 26 OTD per jin, and by early March, it had reached around 30 OTD per jin. Journalist Yang Feng left Taiwan before his article was published in Wenhui Daily on March 4, 1947. Yang Feng said that the price of rice at that time was around 40 OTD per jin. Based on this, it can be inferred that the figure "42.67 OTD per jin" given by scholar Weng Jiaxi might have existed in late February 1947.
@@ 3 Research Methods
Then, we performed a weighted average of different data from the same month. Chen Huai-de believed that official data should be regarded as the main source of trust for data, and its proportion in the weighted average calculation should not be less than 50 percents. After discussion, we finally determined that the weight of the official monthly price report data in the weighted average calculation is 0.7, and the total weight of other data is 0.3. For instance, when estimating the rice price in January 1946, we adopted a weighted average method of 8.84 as mentioned by Weng Jiaxi and 8.2575 as the "official data". We accepted 70 percents of the official data and 30 percents of what Weng Jiaxi said. We stipulated that the error should be within 5 percents, and thus the revised average monthly rice price for that month was 8.4323±0.4 OTD per jin. By analogy, we obtained the monthly average revised value of rice prices in the Taipei market from August 1945 to March 1947.
@@ 4.1 Preliminary Quantitative Analysis
The results are as presented in Figure. 2 and Figure. 3 and Table. 1 and Table. 2. We have obtained the most perfect quantitative results among several attempts. Subsequently, we sorted out the summary table of the regression model, drew the fitting curve graph of the regression model, and uploaded them to the Science Data Bank[9].
@@
Note. DV = monthly average rice price (old Taiwan dollars); time domain: Aug 1945–May 1946; total cases = 7, excluded = 0.
Figure 2: Fitting diagram of the rice price fluctuation model in the first two stages of the grain shortage from August 1945 to May 1946.
@@
Note. DV = monthly average rice price (old Taiwan dollars); time domain: Nov 1946–Mar 1947; total cases = 6, excluded = 0.
Figure 3: Fitting diagram of the rice price fluctuation model in the first two stages of the grain shortage from August 1945 to May 1946.
4.2 Advanced Quantitative Analysis
Firstly, we conducted descriptive statistical analysis on the rice price of the two groups (defined by the Policy dummy variable "policy 0-1 var"). The statistical data of Group 0 and Group 1 are shown in Table.3.
@@ 4.2 Advanced Quantitative Analysis
As shown in Table.8, the significance levels for the Pearson correlation coefficients are presented. All correlation coefficients are significant at the one-tailed level (p < 0.001), confirming the strong statistical relationship between rice prices and the policy variables.
@@
Variable Beta Correlations Collinearity Statistics
@@
Figure 4: Residual statistics chart of the lagging regression model of the Old Taiwan dollar exchange policy impacting rice prices
Figure 5: P-P plot
Figure 6: Standardized residual histogram
We evaluated the performance of the regression model and the satisfaction of its assumptions by analyzing three key charts, as shown in Figure.4,5,6 First, we examined the standardized residual histogram. This figure shows the distribution of the residuals of the regression model. Ideally, the residuals should approach a normal distribution, which is an important prerequisite for the effectiveness of linear regression models. The shape of the histogram in the figure is roughly bell-shaped and symmetrical to zero, which indicates that the residuals of the model are close to a normal distribution, thereby supporting the validity of the model.
@@ 5.1 Four Stages of Rice Price Changes
This study has provided a comprehensive quantitative reassessment of rice price dynamics during the 1945–1947 famine in post-war Taiwan. By analyzing high-frequency rice price data and employing regression models, we identified four distinct stages of rice price changes and explored the correlations between these changes and government policies.
The first stage (August 1945–September 1945) was characterized by stable and slightly declining prices, reflecting the early post-war recovery period. The second stage (September 1945–May 1946) saw a significant increase in rice prices, marking the first peak of the grain shortage. The third stage (May 1946–November 1946) was a period of relative stability and a gradual decline in prices. The fourth stage (November 1946–April 1947) witnessed sharp fluctuations and a sustained high level of rice prices, leading to social unrest.
@@ 5.3 Contributions and Future Research Directions
While this study has provided significant insights, several avenues for future research remain open. Future studies could explore the impact of additional economic variables, such as inflation rates, exchange rates, and agricultural production levels, on rice prices. Moreover, extending the analysis to other regions or time periods could offer a broader perspective on the relationship between policy interventions and food price dynamics. Finally, the application of more advanced econometric techniques, such as panel data models or structural equation models, could enhance the robustness and accuracy of the findings. Of course, not considering other factors in the regression model is to simplify our modeling, to ignore all these factors into "constants", and to focus only on the policy of the Chen Yi government that we have selected. Besides, many studies have confirmed other factors.

### v2

@@ Abstract
Taiwan , The second Sino-Japanese War , Food crisis , 1945–1947 , Rice price
{graphicalabstract}[Uncaptioned image]
{highlights}
The first high-frequency rice price panel for Taiwan from August 1945 to March 1947
Four stages describing the pattern of rice price trend
Highlighting the dominant role of policy systems in post-war food crises
@@ 1.1 Historical Background: The End of the Second Sino-Japanese War
We started narrating the historical background from 1943. On December 1st, 1943, the heads of state of the United States, China, and the United Kingdom jointly issued the Cairo Declaration, which stipulated the ownership of Taiwan after World War II: ”It is their purpose that Japan shall be stripped of all the islands in the Pacific which she has seized or occupied since the beginning of the first World War in 1914, and that all the territories Japan has stolen from the Chinese, such as Manchuria, Formosa [^3] Formosa is another name for Taiwan., and The Pescadores, shall be restored to the Republic of China.”[1]Similarly, the Potsdam Proclamation also stipulated that Taiwan belonged to China after the war.
On August 15th, 1945, Japan surrendered unconditionally, and the sovereignty of Taiwan was legally returned to China. On September 1st of the same year, the Chinese Nanjing National Government established the Taiwan Provincial Garrison Command in Chongqing. On September 3rd, Chiang Kai-shek issued an order: ”Appoint General Chen Yi as the chief official for accepting the surrender of Taiwan and the Penghu Islands.”[2]At 10 a.m. on October 10th, all officers and soldiers of the Forward Command Post, along with local gentry and civilians from Taiwan, gathered at the former Guild Hall [^4] It’s now known as Taipei Zhongshan Hall.in Taipeito hold the first National Day commemoration in Taiwan Province, declaring the return of Taiwan’s sovereignty.
@@
In August 1945, the price of rice once dropped below 2 Old Taiwan dollar(OTD [^5] Old Taiwan dollar (1946-1949), predecessor to today’s New Taiwan dollar, withdrawn at 40,000:1 owing to post-war hyperinflation. We abbreviate the Old Taiwan dollar as OTD.) per jin [^6] Jin is a traditional chinese unit of measurement. Commonly, one jin is equal to 500 grams.. From mid-August to mid-September 1945, the market in Taiwan Province once presented a scene of abundant goods and a steady decline in prices. However, From the second half of September in the 34th year of the Republic of China [^7] September 1945., the price of grain soared. In early 1947, the price of rice in Taipei rose to 14 OTD per jin. Later, under the influence of the gold rush in Shanghai, the price of rice suddenly jumped to over 30 OTD per jin, and even reached more than 40 OTD per jin on the black market.
@@ 2 Literature Review
Some scholars have conducted research on the food shortage during this period. Many scholars have studied the rice price data at some specific time points. Zeng Lei-lei pointed out that the price of rice in Taiwwan rose above 2 OTD per jin on October 10th and soared to 8 OTD/jin by the end of December[4]. Weng Jia-hsi: ”In January 1946, one jin of rice cost only 8.84 OTD.”[5]On page 265 of Weng’s paper, Weng presented a table of the prices of major daily necessities in Taipei City in the early post-war period. Among them, the price of rice was 8.84 OTD/jin in January 1946 and rose to 42.67 OTD/jin in February 1947, with an increase multiple of 4.83[5]. Professor Wang Xiaobo from Taiwan believes that according to the investigation data of some scholars, the price of rice in Taiwan was 6.3 OTD per jin in January 1946 and 32.33 OTD per jin in February 1947. However, we remain skeptical about the data provided by Wang Xiaobo because the gap between the data he provided and those provided by other scholars exceeds 23 percents, and it is even larger than the official data of the Taiwan government. We fully doubted that the data he provided are for other types of rice, not for top-quality rice.
Additionally, we have found many official documentary books that record the food shortage events that occurred in Taiwan during this period. The book ”Taiwan’s Recovery and the Provincial Conditions of the Five Years After Recovery”, edited by Chen Mingzhong and Chen Xingtang and published by Nanjing Publishing House, includes many reports from newspapers and periodicals during the grain shortage incident. We believe that the representative newspaper of these newspaper types is Wenhui Daily. These two individuals are researchers at the Second National Historical Archives of China, as we believe the content of this book is highly credible[6]. Given that this material has not been reprinted since 2008 and it cannot be found in the bookstore, to Science Data Bank[7] we uploaded the scanned copy whose filename is RefBook1.pdf [^8] DOI: 10.57760/sciencedb.32762. In March 1947, several reports in Wenhui Daily mentioned the price of rice in the Taiwanese market at that time. Among them, the reports of the two journalists caught our attention. Journalist Feng Yan’s report mentioned that the price of rice was 13.5 OTD per jin in January 1947, soon rose to 26 OTD per jin, and by early March, it had reached around 30 OTD per jin. Journalist Yang Feng left Taiwan before his article was published in Wenhui Daily on March 4, 1947. Yang Feng said that the price of rice at that time was around 40 OTD per jin. Based on this, it can be inferred that the figure ”42.67 OTD per jin” given by scholar Weng Jiaxi might have existed in late February 1947.
@@ 3 Research Methods
Then, we performed a weighted average of different data from the same month. Chen Huai-de believed that official data should be regarded as the main source of trust for data, and its proportion in the weighted average calculation should not be less than 50 percents. After discussion, we finally determined that the weight of the official monthly price report data in the weighted average calculation is 0.7, and the total weight of other data is 0.3. For instance, when estimating the rice price in January 1946, we adopted a weighted average method of 8.84 as mentioned by Weng Jiaxi and 8.2575 as the ”official data”. We accepted 70 percents of the official data and 30 percents of what Weng Jiaxi said. We stipulated that the error should be within 5 percents, and thus the revised average monthly rice price for that month was 8.4323±0.4 OTD per jin. By analogy, we obtained the monthly average revised value of rice prices in the Taipei market from August 1945 to March 1947.
@@ 4.1 Preliminary Quantitative Analysis
The results are as presented in Figure. 2(a) and Figure. 2(b) and Table. 3 and Table. 3. We have obtained the most perfect quantitative results among several attempts. Subsequently, we sorted out the summary table of the regression model, drew the fitting curve graph of the regression model, and uploaded them to the Science Data Bank[9].
@@
(a) Fitting diagram of the rice price fluctuation model in the first two stages of the grain shortage from August 1945 to May 1946.
(b) Fitting diagram of the rice price fluctuation model in the fourth stage of the grain shortage from November 1946 to March 1947.
Figure 2:
4.2 Advanced Quantitative Analysis
Firstly, we conducted descriptive statistical analysis on the rice price of the two groups (defined by the Policy dummy variable ”policy 0-1 var”). The statistical data of Group 0 and Group 1 are shown in Table.3.
@@ 4.2 Advanced Quantitative Analysis
Variable $\beta$ Correlations Col Stats
@@
(a) Residual statistics chart of the lagging regression model of the Old Taiwan dollar exchange policy impacting rice prices
(b) P-P plot
(c) Standardized residual histogram
Figure 3: Residual analysis
As shown in Table.8, the significance levels for the Pearson correlation coefficients are presented. All correlation coefficients are significant at the one-tailed level (p ¡ 0.001), confirming the strong statistical relationship between rice prices and the policy variables.
We evaluated the performance of the regression model and the satisfaction of its assumptions by analyzing three key charts, as shown in Figure.3(a), Figure.3(b) and Figure.3(c) First, we examined the standardized residual histogram. This figure shows the distribution of the residuals of the regression model. Ideally, the residuals should approach a normal distribution, which is an important prerequisite for the effectiveness of linear regression models. The shape of the histogram in the figure is roughly bell-shaped and symmetrical to zero, which indicates that the residuals of the model are close to a normal distribution, thereby supporting the validity of the model.
@@ 5.1 Four Stages of Rice Price Changes
This study has provided a comprehensive quantitative reassessment of rice price dynamics during the 1945-1947 famine in post-war Taiwan. By analyzing high-frequency rice price data and employing regression models, we identified four distinct stages of rice price changes and explored the correlations between these changes and government policies.
The first stage (August 1945-September 1945) was characterized by stable and slightly declining prices, reflecting the early post-war recovery period. The second stage (September 1945-May 1946) saw a significant increase in rice prices, marking the first peak of the grain shortage. The third stage (May 1946-November 1946) was a period of relative stability and a gradual decline in prices. The fourth stage (November 1946-April 1947) witnessed sharp fluctuations and a sustained high level of rice prices, leading to social unrest.
@@ 5.3 Contributions and Future Research Directions
While this study has provided significant insights, several avenues for future research remain open. Future studies could explore the impact of additional economic variables, such as inflation rates, exchange rates, and agricultural production levels, on rice prices. Moreover, extending the analysis to other regions or time periods could offer a broader perspective on the relationship between policy interventions and food price dynamics. Finally, the application of more advanced econometric techniques, such as panel data models or structural equation models, could enhance the robustness and accuracy of the findings. Of course, not considering other factors in the regression model is to simplify our modeling, to ignore all these factors into ”constants”, and to focus only on the policy of the Chen Yi government that we have selected. Besides, many studies have confirmed other factors.